OpenStudy (anonymous):

Solve the system by graphing. Enter your answer as an ordered pair. y=-x+4 x=y+2

geerky42 (geerky42):

\(\color{red}y = -x+4\\x =\color{red}y+2\longrightarrow x=\color{red}{(-x+4)}+4 = -x+4+4\)